A fruit company guarantees that 92​% of the pineapples it ships will ripen within four days of delivery. Find the probability that at least 10 pineapples in a case of 12 are ripe within four days.

Answer by Boreal(15235)

That is 10, 11, or 12
for 10 it is 12C10*.92^10*.08^2=0.1835
for 11 it is 12*.92^11*.08=0.3837
for 12 it is .92^12=0.3677
sum is 0.9348 probability